OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. x2 - 20x + 100 = 36 HELP!!!!!!!!!!!!!

OpenStudy (btaylor):

First, realize that x^2 - 20x + 100 = (x-10)^2. Then take the positive/negative square roots of 36. Then you have the answer.

OpenStudy (anonymous):

so we are looking to factorise it as \[(x+a)^2 \] if we just expand this ^ for the sake of explaining \[(x+a)^2 = (x+a)(x+a) = x(x+a) + a(x+a) = x^2 + 2ax + a^2\] so \[x^2 +20x+ 100 = x^2 + 2ax +a^2\] so we can see that \[2ax = 20x\] \[2a = 20\] \[a = 10\] so we have \[(x+10)^2 = 36\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

can I get that in a ordered pair ( , )

OpenStudy (anonymous):

\[(x+10)^2 = 36\] \[(x+10) = \pm 6\]\[x = \pm 6 -10\] x= -16 , x = -4

OpenStudy (anonymous):

do you understand the method?